OpenStudy (chevyexpress1962):

What is the quadratic function that is created with roots -3 and 1 and a vertex at (-1, -8)?

OpenStudy (peachpi):

quadratic function with known roots: \[f(x)=a(x-p)(x -q)\] where p and q are the roots. So your function is \[f(x)=a(x+3)(x-1)\] Plug in the vertex and solve for \(a\)

OpenStudy (peachpi):

ANOTHER WAY TO DO THIS. quadratic function with known vertex: \[f(x)=a(x-h)^2+k\] where (h, k) is the vertex So your function is\[f(x)=a(x+1)^2-8\] Plug in one of the roots and solve for \(a\) You'll get the same function with either method.

OpenStudy (chevyexpress1962):

Thank you

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)